OpenStudy (anonymous):

6raised to the x= 20

zepdrix (zepdrix):

solve for x?

OpenStudy (anonymous):

what are the three properties of logs?

zepdrix (zepdrix):

\[a \log(b) = \log(b^a)\] \[\log(a)+\log(b)=\log(a*b)\] other ones? mm i forget off the top of my head.. the first one i listed is the one we need to use here :D

zepdrix (zepdrix):

first step, take the natural log of both sides :D \[6^x=20\rightarrow \ln (6^x)=\ln (20)\]

zepdrix (zepdrix):

see how the first log rule might help us on the left side of the equation?

OpenStudy (anonymous):

for future people!!! the 3 properties of logarithms are the power property,the product property, and the quotient property!. gl on the dba ;)

zepdrix (zepdrix):

yes people!! and remember to drink your applejuice! :O

OpenStudy (anonymous):

that sir, earned you a medal.

zepdrix (zepdrix):

XD